OpenStudy (anonymous):

Please help me, I'm really confused f(x) = {3x^2 - 12x + 2 - 3x + 10 If you could show the work so I can understand it? Thanks!!

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@radar

OpenStudy (anonymous):

@wolfe8

jimthompson5910 (jim_thompson5910):

what are the instructions? do you want to graph this?

OpenStudy (anonymous):

what are you trying to do with it?

OpenStudy (anonymous):

It just says to solve the system @jim_thompson5910 @satellite73

OpenStudy (anonymous):

And then gives that system^^ @jim_thompson5910 @satellite73

OpenStudy (anonymous):

oh ok set them equal and solve for \(x\)

OpenStudy (anonymous):

But what about the exponent of 2?

OpenStudy (anonymous):

start with \[3x^2 - 12x + 2= - 3x + 10\]

OpenStudy (anonymous):

then set equal to zero via \[3x^2-9x-8=0\] then if you are lucky it factors

OpenStudy (anonymous):

Like foil?

OpenStudy (anonymous):

it doesn't factor, use the quadratic formula

OpenStudy (anonymous):

can you post a picture of the exact problem?

OpenStudy (anonymous):

cause it looks like you have function notation there, not a system of equations

OpenStudy (anonymous):

So it'd be \[9 \pm \sqrt{9^2 - 4(3)(-8)}\div(2)(3)\] ??

OpenStudy (anonymous):

yes or \[\frac{9\pm\sqrt{177}}{6}\]

OpenStudy (anonymous):

And how do you know if it is + or -?

OpenStudy (anonymous):

but that is the solution to the system \[y=3x^2 - 12x + 2 \\ y=- 3x + 10 \]

OpenStudy (anonymous):

wait what?

OpenStudy (anonymous):

what you wrote does not look like a system of equations

OpenStudy (anonymous):

you wrote something that looks like this \[f(x) = \left\{\begin{array}{rcc} 3x^2-12x+2 & \ \\ -3x-10& \end{array} \right. \]

OpenStudy (anonymous):

that is why i asked if you could post a picture

OpenStudy (anonymous):